# S-GRPO: Early Exit via Reinforcement Learning in Reasoning Models

### v1

@@ 3.2 Experimental Results
The experimental results in Table 2 demonstrate that our S-GRPO consistently and significantly outperforms existing baselines across five benchmark datasets and four reasoning models. Compared to vanilla reasoning models, S-GRPO achieves an average accuracy improvement of 0.74 to 6.08 percentage points, while reducing the generated sequence length by 35.4% to 61.1%. Our method delivers notable gains both on in-domain mathematical reasoning benchmarks used in training (i.e., GSM8K, AIME-2024, AMC-2023, MATH-500) and on out-of-domain scientific reasoning tasks (i.e., GPQA), providing strong evidence of its effectiveness and robustness. Specifically, when applied to DeepSeek-R1-Distill-Qwen-7B, S-GRPO achieves a 10.3 points increase in accuracy on AMC 2023 while utilizing only 36% of the original reasoning budget. Similarly, it obtains a 0.6 point improvement on AIME 2024 with just 56% of the allocated inference length.
@@
In comparison, DEER reduces reasoning length by only 17% on AIME-2024, yet at the cost of a decline in accuracy. Unlike DEER, whose efficiency is primarily limited to simple tasks, S-GRPO consistently reduces reasoning length across both easy and challenging tasks, while also enhancing reasoning accuracy. Compared to the original GRPO, S-GRPO achieves comparable or even better accuracy while significantly shortening the inference trajectory, indicating that the proposed Serial-Group Generation mechanism does not hinder the model’s exploration capability in reinforcement learning. Moreover, the reward shaping based on serialized intermediate outputs better guides the model toward efficient reasoning.
Compared with the most recent training-based efficient reasoning methods, S-GRPO achieves the best performance on both reasoning length reduction and accuracy. This can be explained by the fact that ConCISE’s off-policy preference optimization forces the model to fit into a new data distribution, and ShorterBetter and RL + Length Penalty overly emphasize length reduction, at the expense of generalization and accuracy-driven reward optimization. In contrast, S-GRPO preserves the integrity of the original reasoning process via a two-stage rollout procedure, ensuring that the model’s pre-existing reasoning abilities are not compromised, which is well-suited as the final stage of post-training.
Figure 3: Performance of DeepSeek-R1-Distill-Qwen-7B and Qwen3-8B without or with S-GRPO training on GSM8K and AIME24 under different generation-length budgets.
@@
We vary the generation-length budget during inference from short to long, and compare the accuracy and actual reasoning length of S-GRPO and vanilla CoT on GSM8K (representing simpler problems) and AIME-2024 (representing more complex problems). The results in Figure 3 show that, across all tested budgets, S-GRPO consistently outperforms vanilla CoT in both accuracy and efficiency with shorter generated sequences, further demonstrating the effectiveness of our method.
Moreover, we observe that under tight length budgets, S-GRPO achieves significantly higher accuracy than vanilla CoT while generating sequences of similar length. In contrast, under larger length budgets, S-GRPO produces significantly shorter reasoning paths with slightly better accuracy. The trends in both accuracy and actual generation length of S-GRPO are smoother compared to vanilla CoT, indicating its greater robustness to variations in length budget. Overall, S-GRPO achieves high accuracy with a much lower length budget, suggesting that our method is capable of generating concise yet accurate reasoning paths.
@@
Method GSM8K AIME24 AMC23 MATH-500 GPQA Overall
@@
S-GRPO 96.1 1292 77.3 8810 95.0 5962 95.2 3166 57.7 5271 84.26 4922
w/o. Decaying (Shortest 1) 95.9 1175 69.6 8721 92.5 4581 94.8 2740 55.7 4734 81.70-2.56 4390-10.8%
w/o. Decaying (All 1) 96.0 2385 74.4 14940 94.7 9000 95.0 5614 54.9 8955 83.00-1.26 8179+66.2%
  - w/o. Serial 95.8 2355 72.7 15154 92.8 8983 94.4 5440 55.8 8819 82.30-1.96 8150+65.6%
Ablation Results.
To verify the effectiveness of each design component in S-GRPO, we conduct ablation studies under three different settings. w/o. Decaying (Shortest 1) denotes the setting where only the shortest correct completion sampled from the two-time rollouts is assigned a reward of 1, while all other samples receive a reward of 0. w/o. Decaying (All 1) refers to the configuration in which all correctly answered completions sampled via two-time rollouts are assigned a reward of 1, while incorrect ones receive a reward of 0. w/o. Serial indicates that, based on the w/o. Decaying (All 1) setup, we further remove the Serial-Group Generation mechanism.
The results in Table 2 indicate that assigning rewards only to the shortest correct completion imposes an overly strict constraint. Although this setting leads to further reductions in reasoning length, it comes at the cost of reasoning accuracy. Removing the design that assigns higher rewards to shorter outputs, i.e., w/o. Decaying (All 1), still results in lengthy reasoning. We attribute this to the fact that longer reasoning sequences are more likely to lead to correct answers, as suggested by test-time scaling theory, causing the model to ultimately converge toward exploring longer reasoning paths. When the Serial-Group Generation mechanism is removed, our method degenerates to GRPO, achieving performance comparable to w/o. Decaying (All 1) in both accuracy and reasoning length. This demonstrates that Serial-Group Generation, as an essential component of S-GRPO, does not compromise the model’s exploration capability in reinforcement learning.
@@
Figure 5 in the Appendix visually illustrates how S-GRPO samples completions that exit early at different positions during the second rollout, along with the corresponding reward assignments. For the first early-exit position, where the model produces an incorrect intermediate answer, we set the reward to zero. For subsequent exits that yield correct answers, we apply a decaying positive reward scheme, where earlier exits are associated with higher reward values. This design incentivizes the model to discover reasoning paths that are both accurate and succinct.
To more intuitively illustrate the effectiveness of our approach, we present a representative example in Figure 4. The generated contents on the left and right sides of the figure illustrate the reasoning processes of S-GRPO and vanilla CoT, respectively. Although both methods yield the correct final answer, S-GRPO achieves this with less than half of the allocated reasoning budget, demonstrating its ability to effectively mitigate the overthinking problem chen2025think23overthinkingo1like . The central portion of the figure displays an early exit obtained by directly truncating the original model’s reasoning process at the same thinking budget as our method. The reasoning model fails to reach the correct conclusion based on the available partial reasoning. This indicates that S-GRPO more accurately identifies the correct solution path and guides the model toward concise yet accurate reasoning, inherently avoiding the underthinking issue where each reasoning path is only explored superficially wang2025thoughtsplaceunderthinkingo1like .
@@ 4 Related Work
Reinforcement Learning for Reasoning
Rule-based outcome reward RL deepseekai2025deepseekr1incentivizingreasoningcapability ; kimiteam2025kimik15scalingreinforcement ; gao2024designingeffectiverlreward ; lambert2025tulu3pushingfrontiers ; zeng2025simplerlzooinvestigatingtamingzero ; wen2025lightr1curriculumsftdpo ; song2025fastcurlcurriculumreinforcementlearning has emerged as the mainstream approach for post-training optimization of reasoning models. This paradigm simplifies reward design by employing binary 0/1 rewards determined through rule-based correctness evaluation, eliminating the need for separate reward models as required in original GRPO shao2024deepseekmathpushinglimitsmathematical ; deepseekai2025deepseekr1incentivizingreasoningcapability ; schulman2017proximalpolicyoptimizationalgorithms implementations, thereby substantially reducing memory and computational overhead during RL training. Recent advances have introduced numerous RL algorithm variants focusing on training efficiency yu2025dapoopensourcellmreinforcement ; liu2025understandingr1zeroliketrainingcritical ; deepcoder2025 ; zhang2025srpocrossdomainimplementationlargescale , value function optimization kazemnejad2024vineppounlockingrlpotential ; yuan2025whatspposcollapselongcot ; yue2025vapoefficientreliablereinforcement , and other aspects. However, these developments preserve GRPO’s fundamental mechanism that samples several CoTs in parallel for a group, consequently overlooking optimization opportunities during sequential thought generation within a single CoT. Our proposed optimized version of GRPO, S-GRPO, maintains the rule-based outcome reward framework while enabling intermediate reasoning process rewards by sampling multiple early-exit thought chains in serial for one CoT. We allow existing reasoning models to effectively mitigate overthinking and boost inference efficiency.
@@
Training-based methods primarily supervised fine-tuning models with variable-length CoT data yu2024distilling ; kang2025c3ot ; xia2025tokenskip ; ma2025cot ; munkhbat2025self ; liu2024can ; han2024token , or training with length-rewards RL team2025kimi ; luo2025o1 ; aggarwal2025l1 ; arora2025training ; yeo2025demystifying ; shen2025dast ; qu2025optimizing ; cui2025stepwise . Recently, ConCISE qiao2025conciseconfidenceguidedcompressionstepbystep constructs concise CoT data by inserting prompt tokens and employing early-exit during inference, then enhances the model’s reasoning conciseness through DPO rafailov2024directpreferenceoptimizationlanguage ; meng2024simposimplepreferenceoptimization . arora2025traininglanguagemodelsreason and yi2025shorterbetterguidingreasoningmodels upweight rewards for the shorter CoTs, having correct answers, in parallel-sampled reasoning chains. However, they inherit GRPO’s parallel sampling paradigm that neglects the reward function’s attention to the intermediate serial reasoning processes.
@@ 5 Conclusion
This paper proposes Serial-Group Decaying-Reward Policy Optimization, which innovatively leverages rule-based outcome rewards to regulate intermediate reasoning processes, reinforcing the model to generate high-quality thoughts earlier and exit early when the generated thought is sufficient. Experiments demonstrate that S-GRPO consistently achieves the best efficiency-accuracy synergistic improvement, significantly outperforming existing efficient reasoning methods across five benchmark datasets and four reasoning models. We recommend deploying S-GRPO as the final optimization stage in post-training pipelines.

### v2

@@ 3.2 Experimental Results
The experimental results in Table 1 demonstrate that S-GRPO consistently outperforms existing baselines across five benchmark datasets and four reasoning models, achieving significant improvements. Compared to vanilla reasoning models, S-GRPO achieves an average accuracy (absolute) improvement of 0.72% to 6.08%, while reducing the generated sequence length by 35.4% to 61.1%. S-GRPO achieves notable improvements on in-domain mathematical reasoning benchmarks (e.g., GSM8K, AIME 2024, AMC 2023, MATH-500) and out-of-domain scientific reasoning tasks (e.g., GPQA), demonstrating its effectiveness and robustness. Specifically, when applied to DeepSeek-R1-Distill-Qwen-7B, S-GRPO achieves a 10.3-point increase in accuracy on AMC 2023 while utilizing only 36% of the original reasoning budget. Similarly, it obtains a 0.6-point improvement on AIME 2024 with just 56% of the original inference length.
@@
We compare S-GRPO with several approaches designed for efficient reasoning. For example, DEER reduces reasoning length by only 17% on AIME 2024 and sacrifices accuracy, and its effectiveness is mainly limited to simple tasks. In contrast, S-GRPO consistently reduces reasoning length across both simple and complex tasks while simultaneously improving reasoning accuracy. Compared to the original GRPO, S-GRPO achieves comparable or even better accuracy while significantly shortening the inference trajectory, indicating that the proposed Serial-Group Generation mechanism does not hinder the model’s exploration capability in reinforcement learning. Moreover, the reward shaping based on serialized intermediate outputs more effectively guides the model toward efficient reasoning.
Compared with the most recent training-based efficient reasoning methods, S-GRPO achieves the best in both reasoning length reduction and accuracy. This can be explained by the fact that ConCISE’s off-policy optimization forces the model to fit into a new data distribution, and ShorterBetter and RL + Length Penalty overly emphasize length reduction, at the expense of generalization and accuracy-driven optimization. In contrast, S-GRPO preserves the integrity of the original reasoning process via a two-stage rollout procedure, ensuring that the model’s pre-existing reasoning abilities are not compromised, which is well-suited as the final stage of post-training.
Figure 3: Performance of DeepSeek-R1-Distill-Qwen-7B and Qwen3-8B without or with S-GRPO training on GSM8K and AIME 2024 under different generation-length budgets.
@@
We vary the generation-length budget during inference from short to long, and compare the change in accuracy and actual reasoning length of S-GRPO and vanilla CoT on GSM8K (representing simple problems) and AIME 2024 (representing complex problems). The results in Figure 3 show that, across all budgets tested, S-GRPO consistently achieves higher accuracy while generating shorter sequences compared to vanilla CoT, further highlighting the effectiveness of our method.
Moreover, we observe that under tight length budgets, S-GRPO achieves significantly higher accuracy than vanilla CoT while generating sequences with comparable length. Differently, under larger length budgets, S-GRPO produces significantly shorter reasoning paths with slightly better accuracy, compared with vanilla CoT. The trends in both accuracy and actual generation length of S-GRPO are smoother than vanilla CoT, indicating its greater robustness to variations in length budget. Overall, S-GRPO achieves high accuracy under a low length budget, suggesting that our method is capable of generating concise yet accurate reasoning paths.
Ablation results.
To verify the effectiveness of each design component in S-GRPO, we conduct ablation studies under three different settings. w/o. Decaying (Shortest 1) denotes the setting where only the shortest correct response in the serial group is assigned a reward of 1, while all other responses receive a reward of 0. w/o. Decaying (All 1) refers to the configuration where all correct responses in the serial group are assigned a reward of 1, while incorrect ones receive a reward of 0. w/o. Serial indicates that we further remove the Serial-Group Generation mechanism based on the w/o. Decaying (All 1) configuration
@@
Method GSM8K AIME 2024 AMC 2023 MATH-500 GPQA Overall
@@
S-GRPO 96.1 1,292 77.3 8,810 95.0 5,962 95.2 3,166 57.7 5,271 84.26 4,922
w/o. Decaying (Shortest 1) 95.9 1,175 69.6 8,721 92.5 4,581 94.8 2,740 55.7 4,734 81.70-2.56 4,390-10.8%
w/o. Decaying (All 1) 96.0 2,385 74.4 14,940 94.7 9,000 95.0 5,614 54.9 8,955 83.00-1.26 8,179+66.2%
  - w/o. Serial 95.8 2,355 72.7 15,154 92.8 8,983 94.4 5,440 55.8 8,819 82.30-1.96 8,150+65.6%
The results in Table 2 indicate that rewarding only the shortest correct response imposes an overly strict constraint. Although this setting leads to further reductions in reasoning length, it comes at the cost of accuracy. And removing the design that assigns higher rewards to shorter outputs, i.e., w/o. Decaying (All 1), results in lengthy reasoning. This is because long correct answers also receive high rewards, and the model does not shift in favor of generating short CoTs. When the Serial-Group Generation mechanism is removed, our method degenerates to GRPO, achieving performance comparable to w/o. Decaying (All 1) in both accuracy and reasoning length. This demonstrates that Serial-Group Generation, as an essential component of S-GRPO, does not compromise the model’s exploration capability in reinforcement learning.
@@
Figure 5 in the Appendix visually illustrates how S-GRPO generates answers at different early-exit positions and also shows the decaying reward assignments. For the first early-exit position, where the model produces an incorrect intermediate answer, we set the reward to 0. For subsequent exits that yield correct answers, we apply a decaying positive reward scheme, where earlier exits are associated with higher reward values. This design incentivizes the model to discover reasoning paths that are both accurate and succinct.
To more intuitively illustrate the effectiveness of our approach, a representative example is shown in Figure 4. The left and right sides of the figure compare the reasoning processes of vanilla CoT and S-GRPO. Although both methods yield the correct final answer, S-GRPO achieves this using less than half of the reasoning budget, showcasing its effectiveness in mitigating the overthinking problem chen2025think23overthinkingo1like . The central portion of the figure displays an early exit obtained by directly truncating the vanilla CoT’s reasoning process using the same token budget as S-GRPO. The reasoning model fails to reach the correct conclusion based on the available partial reasoning. This demonstrates that S-GRPO effectively identifies the correct solution path and guides the model to concise and accurate reasoning, while inherently avoiding the underthinking issue of superficial exploration wang2025thoughtsplaceunderthinkingo1like .
@@ 4 Related Work
Reinforcement learning for reasoning.
Rule-based outcome reward RL deepseekai2025deepseekr1incentivizingreasoningcapability ; kimiteam2025kimik15scalingreinforcement ; gao2024designingeffectiverlreward ; lambert2025tulu3pushingfrontiers ; zeng2025simplerlzooinvestigatingtamingzero ; wen2025lightr1curriculumsftdpo ; song2025fastcurlcurriculumreinforcementlearning has emerged as the mainstream approach for post-training optimization of reasoning models. This paradigm simplifies reward design by employing binary 0/1 rewards determined through rule-based correctness evaluation, eliminating the need for separate reward models as required in original GRPO shao2024deepseekmathpushinglimitsmathematical ; deepseekai2025deepseekr1incentivizingreasoningcapability ; schulman2017proximalpolicyoptimizationalgorithms implementations, thereby substantially reducing memory and computational overhead during RL training. Recent advances have introduced numerous RL algorithm variants focusing on training efficiency yu2025dapoopensourcellmreinforcement ; liu2025understandingr1zeroliketrainingcritical ; zhang2025srpocrossdomainimplementationlargescale , value function optimization kazemnejad2024vineppounlockingrlpotential ; yuan2025whatspposcollapselongcot ; yue2025vapoefficientreliablereinforcement , and other aspects. However, these developments preserve GRPO’s fundamental mechanism that samples several CoTs in parallel for a group, consequently overlooking optimization opportunities during sequential thought generation within a single CoT. Our proposed optimized version of GRPO, S-GRPO, maintains the rule-based outcome reward framework while enabling intermediate reasoning process rewards by sampling multiple early-exit thought chains in serial for one CoT. We allow existing reasoning models to effectively mitigate overthinking and boost inference efficiency.
@@
Training-based methods primarily supervised fine-tuning models with variable-length CoT data yu2024distilling ; kang2025c3ot ; xia2025tokenskip ; ma2025cot ; munkhbat2025self ; liu2024can ; han2024token , or training with length-rewards RL team2025kimi ; luo2025o1 ; aggarwal2025l1 ; arora2025training ; yeo2025demystifying ; shen2025dast ; qu2025optimizing ; cui2025stepwise . Recently, ConCISE qiao2025conciseconfidenceguidedcompressionstepbystep constructs concise CoT data by inserting prompt tokens and employing early-exit during inference, then enhances the model’s reasoning conciseness through SFT/SimPO rafailov2024directpreferenceoptimizationlanguage ; meng2024simposimplepreferenceoptimization . arora2025traininglanguagemodelsreason and yi2025shorterbetterguidingreasoningmodels upweight rewards for the shorter CoTs, having correct answers, in parallel-sampled reasoning chains. However, they inherit GRPO’s parallel sampling paradigm that neglects the reward function’s attention to the intermediate serial reasoning processes.
@@ 5 Conclusion
This paper introduces Serial-Group Decaying-Reward Policy Optimization (S-GRPO), which innovatively leverages rule-based outcome rewards to regulate intermediate reasoning processes. By incentivizing the LLM to produce high-quality thoughts earlier and exit promptly when the generated thought is sufficient, S-GRPO enhances reasoning efficiency and maintains accuracy. Empirical evaluations demonstrate that S-GRPO achieves optimal synergy between efficiency and accuracy, significantly outperforming existing efficient reasoning methods. Specifically, S-GRPO demonstrates compatibility with state-of-the-art reasoning models, including Qwen3 and Deepseek-distill, achieving a 35.4$\sim$61.1reduction in sequence length while improving accuracy (absolute) by 0.72$\sim$6.08across five benchmark datasets: GSM8K, AIME 2024, AMC 2023, MATH-500, and GPQA Diamond. S-GRPO is well-suited for deployment as the final optimization stage in post-training pipelines, offering a practical solution for enhancing reasoning efficiency.